OpenStudy (anonymous):

PLEASE help me practice solving systems 3 equations, 3 variables, 3 answers. PLEASE help me practice solving systems?? I get to re-take a systems quiz tomorrow and I want to do better on it. PLEASE HELP will REWARD AN MEDAL..

OpenStudy (anonymous):

I don't have a particular problem. I'm wanting help solving sysems? with 3 equations... I don't have a particular problem?

OpenStudy (jdoe0001):

http://www.youtube.com/watch?v=aPKz7011lAA maybe ?

OpenStudy (anonymous):

doesn't help me

OpenStudy (jdoe0001):

hmm... is pretty thorough IMO

OpenStudy (jdoe0001):

lemme cook one up then

OpenStudy (dumbcow):

in general use elimination to eliminate a variable and reduce it to a system of 2 equations , then solve that system , then get 3rd variable by back substituting

OpenStudy (jdoe0001):

\(\large \begin{array}{llll} 2x-2z+3y=-24\\ 5z+3y=27\\ 5y-3z=-47 \end{array}\)

OpenStudy (jdoe0001):

hmm I have a typo

OpenStudy (jdoe0001):

\(\large \begin{array}{llll} 2x-2z+3y=-24\\ 5z+3x=27\\ 5y-3z=-47 \end{array}\)

OpenStudy (jdoe0001):

so first off... we want to order all variables, just like in a 2 variable set, in a column notice the 2nd and 3rd rows, they seem to be missing a variable, thus we PAD it with 0 since 0x = " "

OpenStudy (jdoe0001):

so we end up with then \(\large \begin{array}{llll} 2x+3y-2z=-24\\ 3x+0y+5z=27\\ 0x+5y-3z=-47 \end{array}\)

OpenStudy (jdoe0001):

so as the video showed, we start by grabbing a pair first from the 3 available then we cancel out a variable, say "z"

OpenStudy (jdoe0001):

2x+3y-2z=-24 <---- let us pick this one 3x+0y+5z=27 <--- and this one for now 0x+5y-3z=-47

OpenStudy (jdoe0001):

\(\large {\begin{array}{llll} 2x+3y-2z=-24 & \times 5\implies &10x+15y-10z=-120\\ 3x+0y+5z=27& \times 2\implies &6x+0y+10z=54\\ \hline\\ &&16x+15y+0z=-66 \end{array}\\ \quad \\ \color{blue}{16x+15y=-66}} \)

OpenStudy (jdoe0001):

so by using the elimination method, we ended up with the one above, our 1st 2 variable function

OpenStudy (anonymous):

@jdoe0001 I understand it all up until ur entire last message I don't understand ur last message sir/mam????

OpenStudy (jdoe0001):

ok... .so... what part confused you?

OpenStudy (anonymous):

the right side how u transformed the problem to the new two problems??

OpenStudy (jdoe0001):

\(\large \begin{array}{llll} 2x+3y-2z=-24 & \huge{\color{red}{\times 5}}\implies &10x+15y-10z=-120\\ 3x+0y+5z=27& \huge {\color{red}{\times 2}}\implies &6x+0y+10z=54\\ \hline\\ &&16x+15y+0z=-66 \end{array}\\ \quad \\ 16x+15y=-66 \)

OpenStudy (jdoe0001):

see it now?

OpenStudy (jdoe0001):

the idea being, I'd end up with a "z" value atop, EQUAL BUT DIFFERENT SIGN as the one at the bottom

OpenStudy (anonymous):

yupp now I understand I undertand it now thanks

OpenStudy (jdoe0001):

so let us pick another 2 functions from the available 3 2x+3y-2z=-24 <--- this one again 3x+0y+5z=27 0x+5y-3z=-47 <---- and this one this time

OpenStudy (anonymous):

I undertand it now I watched the videos thanks so much for ur help... and u explaining it helped too. I think get it now

OpenStudy (jdoe0001):

\(\large {\begin{array}{llll} 2x+3y-2z=-24 & \color{red}{\times -3}\implies &-6x-9y+6z=72\\ 0x+5y-3z=-47& \color{red}{\times 2}\implies &0x+10y-6z=-94\\ \hline\\ &&-6x+1y+0z=-22 \end{array}\\ \quad \\ \color{blue}{-6x+y=-22}}\)

OpenStudy (jdoe0001):

so there, then you use the 2 resulting 2 variables functions and solve for either then get the other and then plug them back in any of the originals and to get "z" :)

OpenStudy (anonymous):

(-= THANKS SO MUCH..

OpenStudy (jdoe0001):

yw